Four identical spheres each of radius 10 cm and mass
1 kg are placed on a horizontal surface touching one
another so that their centres are located at the corners of square of side 20cm. find distance of their centre of mass from centre of each sphere
 

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Given four identical mass placed at the corner of a square, so one can directly say the center of mass will be at the center of the square. 
Alternatively, Assume one of them at (0,0) the rest of them will be at (a, 0)  (\sqrt{2}a,\sqrt{2} a ),and (0,a)  where a is the side =10cm

So X=(m1x1+m2x2+m3x3+m4x4)/4m =(0+a*10+\sqrt{2} a *10 + 0) /4*10 =(1+\sqrt{2})/4
similiary Y= 1+\sqrt{2})/4